OpenStudy (anonymous):

find an angle between 0 and 2pi coterminal with 35pi/6

OpenStudy (tkhunny):

2pi is the magic hint. 35pi/6 - 2pi = 35pi/6 - 12pi/6 = 23pi/6 -- Not yet. 23pi/6 - 12pi/6 = 11pi/6 -- Done!

OpenStudy (precal):